# Generative AI for Predicting 2D and 3D Wildfire Spread: Beyond Physics-Based Models and Traditional Deep Learning

### v1

@@ Abstract
Wildfires continue to inflict devastating human, environmental, and economic losses globally, as tragically exemplified by the 2025 Los Angeles wildfire and the urgent demand for more effective response strategies. While physics-based and deep learning models have advanced wildfire simulation, they face critical limitations in predicting and visualizing multimodal fire spread in real time, particularly in both 2D and 3D spatial domains using dynamically updated GIS data. These limitations hinder timely emergency response, infrastructure protection, and community safety. Generative AI has recently emerged as a transformative approach across research and industry. Models such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), Transformers, and diffusion-based architectures offer distinct advantages over traditional methods, including the integration of multimodal data, generation of diverse scenarios under uncertainty, and improved modeling of wildfire dynamics across spatial and temporal scales. This position paper advocates for the adoption of generative AI as a foundational framework for wildfire prediction. We explore how such models can enhance 2D fire spread forecasting and enable more realistic, scalable 3D simulations. Additionally, we employ a novel human-AI collaboration framework using large language models (LLMs) for automated knowledge extraction, literature synthesis, and bibliometric mapping. Looking ahead, we identify five key visions for integrating generative AI into wildfire management: multimodal approaches, AI foundation models, conversational AI systems, edge-computing-based scenario generation, and cognitive digital twins. We also address three major challenges accompanying these opportunities and propose potential solutions to support their implementation.
@@ 1 Introduction
Wildfires and bushfires have emerged as one of the most destructive natural disasters of the 21st century, leaving a devastating trail across natural ecosystems, agricultural lands, and densely populated urban regions (Bowman et al., 2020; Doerr and Santín, 2016). These fast-moving fires not only cause immense structural damage and loss of human life but also result in widespread economic disruption and long-term environmental degradation. The release of smoke, fine particulate matter, and toxic gases during large-scale bushfires contributes significantly to air pollution, impacting public health and exacerbating climate change in cities across the globe (Johnston et al., 2012; Urbanski, 2014). A stark example of this escalating crisis was the 2025 Los Angeles wildfire season, during which the Palisades and Eaton wildfires tore through Southern California. Together, they caused an estimated \$250 billion in economic losses, destroyed thousands of homes, and displaced entire communities, inflicting deep physical and emotional suffering on affected populations (EPA, 2025; Gazette, 2025). These events underscore the urgent need for accurate and timely wildfire forecasting systems. In particular, real-time bushfire propagation simulation-capable of predicting fire spread pathways and identifying at-risk zones-plays a critical role in supporting firefighting operations, emergency evacuation planning, and long-term fire hazard mitigation strategies (Baptiste Filippi et al., 2009; Sullivan, 2009c; Wang and Zlatanova, 2019; Wang et al., 2014).
@@
A variety of wildfire propagation models have been developed, each rooted in distinct computational approaches. Physics-based models (e.g., FARSITE, SPARK, Prometheus) simulate fire spread using physical laws of combustion, heat transfer, fuel conditions, and wind dynamics (Finney, 1998; Mell et al., 2007; Miller et al., 2015), offering detailed outputs but requiring extensive computation and environmental inputs—often limiting their use in real-time scenarios (Dipierro et al., 2024). Empirical models, like McArthur’s Fire Danger Index, rely on historical fire behavior to produce rapid, region-specific predictions, but struggle with generalizability across diverse ecosystems (Noble et al., 1980). Traditional machine learning models (e.g., decision trees, support vector machines) have been used for fire ignition and risk classification (Jain et al., 2020), though their capacity to simulate dynamic fire progression is limited. In contrast, deep learning models have recently gained momentum for 2D fire spread forecasting, due to their ability to capture complex spatiotemporal relationships in fire behavior (Andrianarivony and Akhloufi, 2024; Shadrin et al., 2024). Architectures such as CNNs, RNNs, and U-Net variants have been applied to wildfire segmentation and short-term spread prediction, leveraging remote sensing, meteorological, and topographic data (Ghali and Akhloufi, 2023; Vargas, 2025). These deep learning models (e.g., CNNs and RNNs) often outperform traditional physics-based simulators in runtime, scalability, and—in data-rich environments—predictive accuracy, as they learn complex patterns and variability directly from empirical observations rather than relying on explicit physical formulations (Radke et al., 2019). However, deep learning-based fire models still face key limitations that prevent them from enabling more advanced fire prediction capabilities with higher spatial and temporal resolution in real time. Most models remain constrained to 2D spatial domains (Andrianarivony and Akhloufi, 2024) and lack integration and real-time augmentation of multimodal inputs, such as meteorological, topographic, and vegetation (fuel) data (Radke et al., 2019; Abdollahi and Yebra, 2025), which often change dynamically as fire propagates across the landscape. In addition, many existing models cannot simulate vertical fire dynamics, limiting their effectiveness during fast-evolving wildfire events (Xu et al., 2024f).
Building on this direction, emerging generative AI models—such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), Transformers, and diffusion models—offer transformative advantages over earlier deep learning architectures like CNNs and RNNs, particularly in the context of complex environmental modeling tasks such as fire propagation. Unlike traditional models that are often constrained by localized spatial features, limited sequential memory, and sparse or missing data, generative AI models can learn expressive latent representations, capture complex spatiotemporal dependencies, and generate high-resolution outputs from sparse or heterogeneous inputs (Kingma et al., 2013; Goodfellow et al., 2014; Vaswani et al., 2017; Ho et al., 2020). For instance, Transformers enable global attention across long time-series and spatial domains, making them particularly effective in fusing multimodal datasets such as remote sensing imagery, topographic data, real-time meteorological feeds, and vegetation maps (Dosovitskiy et al., 2020; Tay et al., 2022). Diffusion models and GANs have demonstrated the ability to synthesize highly realistic spatial patterns, which can enhance the fidelity of fire spread predictions even under uncertain or rapidly changing conditions (Rombach et al., 2022; Saharia et al., 2022). These strengths position generative models as a promising foundation for real-time, high-resolution bushfire forecasting systems capable of adapting to dynamic environments with greater accuracy and robustness.
Motivated by this potential, the objective of this position paper is to investigate how the convergence of generative AI techniques and emerging multimodal datasets can enable the development of next-generation bushfire prediction systems. Specifically, we explore how these technologies can facilitate real-time data fusion, enable high-fidelity spatial modeling, and enhance forecasting accuracy—capabilities that are essential for improving situational awareness and operational decision-making during fire emergencies (Dilo and Zlatanova, 2011). By delivering timely and reliable predictions, such systems hold the potential to significantly strengthen firefighting coordination, evacuation planning, and strategic fire management. To support this investigation, we first conduct an overview review of existing deep learning approaches applied to fire propagation prediction, focusing on their strengths and limitations. Distinct from prior studies, we adopt a representation learning-driven methodology that leverages Large Language Models (LLMs) and Retrieval-Augmented Generation (RAG) to automate knowledge extraction from over 150 research articles retrieved from academic databases such as IEEE Xplore and Scopus using specialized queries. We then examine recent advancements in generative AI—including VAEs, GANs, Transformers, and diffusion models—that have been used to simulate and forecast fire dynamics, highlighting their advantages over conventional deep learning methods. Finally, we extend this analysis by evaluating how these emerging models can be repurposed or adapted to enable enhanced predictive accuracy, support data augmentation, and facilitate multimodal fire dynamics simulation and real-time fire behavior modeling based on dynamic input data. Looking ahead, we outline five strategic directions to guide this integration: (1) the development of multimodal approaches, (2) the deployment of AI foundation models, (3) the implementation of conversational AI systems, (4) edge-computing-based generation of wildfire scenarios, and (5) the advancement of cognitive digital twins. Alongside these visions, we examine three key challenges that accompany this paradigm shift and propose potential solutions to address them.
@@ 2 Previous Reviews in Fire Spread Management
Several previous studies have conducted comprehensive literature reviews on predicting bushfire spread using simulation techniques, machine learning, and deep learning. This section begins by summarizing these reviews, highlighting how both simulation models and deep learning approaches have been applied to fire propagation prediction. The overview of previous review articles aims to provide a comprehensive understanding of the existing AI applications in wildfire spread prediction and to identify gaps where generative AI models can be leveraged to address challenges faced by traditional simulation and deep learning approaches.
@@ 2.1 Wildfire Simulation Models and Traditional Machine Learning
Wildfire simulation models are traditional, well-established tools that simulate and predict bushfire spread by solving physical and empirical equations based on fuel, weather, and terrain conditions—originating from foundational models like Rothermel’s in the 1970s (Rothermel, 1972; Finney et al., 2012), and they have long served as the backbone of operational fire management and planning systems. Based on their underlying rationale, these models can be classified into several categories, as presented in Table 1, and are summarized through multiple previous studies (Sullivan, 2009b, d, a; Finney et al., 2012).
@@
On the application level, a recent study conducts a comprehensive review of empirical and dynamic wildfire simulation models, focusing on their applications in predicting bushfire and wildfire spread across Australia (Singh et al., 2025). It critically evaluates a suite of simulation systems including PHOENIX Rapidfire, SPARK, AUSTRALIS, REDEYE, IGNITE, and SiroFire, each leveraging distinct modeling techniques to simulate fire behavior, together with a variety of traditional machine learning and deep learning methods for wildfire prediction. PHOENIX Rapidfire, a deterministic simulator, uses a fire characterization model based on Huygen’s principle to predict spread, flame height, and intensity. While it excels in speed and ease of use, it is limited by reliance on predefined behavior models and underestimation of irregular fire shapes (Pugnet et al., 2013; Richards, 1995). SPARK provides flexible, modular simulation driven by user-defined spread models and integrates well with GIS platforms, though its high computational demands pose practical constraints. AUSTRALIS employs a discrete-event simulation based on empirically derived rate-of-spread models; despite its efficiency, it underperforms in severe fire conditions (Miller et al., 2015). REDEYE and IGNITE integrate geospatial and real-time data for risk prediction and hotspot mapping, respectively, yet their effectiveness is limited by data compatibility and platform-specific requirements. SiroFire models dynamic weather components and supports strategic planning but lacks adaptability for diverse fuel types (Singh et al., 2025). Collectively, these simulators provide valuable predictive capabilities; however, each exhibits distinct limitations, ranging from computational inefficiencies and limited model adaptability to reduced accuracy under extreme conditions. This review highlights the need for hybrid modeling approaches that integrate traditional simulation techniques with machine learning and real-time data to enhance predictive accuracy, reduce simulation runtime, and strengthen operational resilience.
@@ 2.2 Deep Learning in Wildfire Prediction
In the following subsection, we leverage a LLM to generate a literature summary and bibliographic visualizations that highlight the evolving trends in deep learning applications over time. Jain et al. (2020) presents a comprehensive scoping review of machine learning (ML) applications in wildfire science and management, analyzing 300 studies published up to the end of 2019. ML usage is categorized across six core domains: fuel characterization, fire detection, climate interactions, risk prediction, fire behavior, and post-fire effects. Among modern approaches, deep learning (DL) algorithms have shown strong potential for modeling wildfire spread, particularly due to their capacity to handle high-dimensional and multivariate data. Convolutional Neural Networks (CNNs) are widely used for spatial fire detection and smoke recognition from imagery, while Long Short-Term Memory (LSTM) networks incorporate temporal dynamics to model fire growth. Deep Neural Networks (DNNs) have also been applied to map burned areas and forecast fire spread, often outperforming traditional models when sufficient annotated data are available. Despite their promise, the review emphasizes the importance of domain expertise and high-quality data for meaningful DL integration in operational forecasting.
A recent study systematically reviews deep learning (DL) applications in forest fire prediction, analyzing 55 key publications from 2017 to 2024 (Mambile et al., 2024). The review emphasizes the transformative role of deep learning in capturing complex spatiotemporal patterns associated with fire ignition and spread, offering significant advantages over traditional machine learning approaches. CNNs are effectively used for analyzing satellite imagery to detect fire-prone areas and assess post-fire damage. LSTM and Gated Recurrent Unit (GRU) networks model time-series data to forecast fire progression based on historical environmental conditions. Emerging generative AI models such as GANs have been used for synthetic data generation, while Multilayer Perceptrons (MLPs) and U-Net architectures support fire risk estimation and boundary segmentation. Despite promising results, DL models face challenges such as data heterogeneity, limited inclusion of human activity factors, generalization across geographies, and the scarcity of high-quality labeled datasets. These limitations highlight the need for integrating diverse data sources and establishing standardized evaluation protocols to ensure model reliability in real-world wildfire prediction scenarios.
@@ 2.3 Limitations of the Existing Deep Learning Applications in Wildfire Prediction
Traditional models such as CNNs and RNNs typically produce deterministic outputs and struggle to quantify prediction uncertainty, an essential capability for wildfire management and emergency response applications involving stochastic environmental processes, such as abrupt changes in wind speed and directions (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017).
@@
Recurrent Neural Networks (RNNs) and Deep Neural Networks (DNNs) often struggle with vanishing gradients and limited temporal memory, which undermines their ability to capture long-range dependencies. This limitation makes them less effective in modeling the temporal progression of wildfires and in representing the long-term variability of underlying environmental processes such as fuel accumulation, climate patterns, and vegetation dynamics (Hochreiter et al., 2001; Chen et al., 2024).
@@
Traditional deep learning architectures are not inherently designed to fuse multimodal data sources, such as 2D GIS data (e.g., satellite imagery, fuel load maps, and meteorological data) and 3D point clouds (e.g., digital terrain models) (Li et al., 2022). As a result, generating multimodal fire spread prediction outputs across different dimensions—from 1D time series to 3D spatial representations—within a unified deep learning framework remains a significant challenge.
@@
Traditional deep learning models, such as CNNs and RNNs, typically achieve strong performance only when trained on abundant, high-quality labeled data. They depend heavily on large volumes of annotated samples, which are often scarce in wildfire prediction tasks due to the rare, unpredictable, and spatially heterogeneous nature of fire events (Ghali and Akhloufi, 2023; Xu et al., 2024f). However, many of these models lack the capability to perform data augmentation or generate synthetic training samples, limiting their predictive accuracy and generalizability in data-sparse or unseen regions. This limitation poses a significant challenge for developing robust wildfire.
@@
Environmental datasets frequently contain missing or incomplete information caused by sensor malfunctions, occlusion from cloud cover in satellite imagery, or data transmission failures (Shen et al., 2015). Such data gaps hinder accurate modeling and prediction. Traditional deep learning models often assume complete input data or rely on simplistic imputation methods that fail to capture the underlying spatiotemporal dependencies critical for wildfire dynamics. These limitations reduce model robustness and prediction accuracy in real-world scenarios where data is inherently noisy or sparse (Li et al., 2020).
@@
Deep learning models such as CNNs and RNNs often operate as ”black boxes,” providing limited insight into how predictions are made and offering little transparency or trustworthiness (Rudin, 2019; Doshi-Velez and Kim, 2017).

### v2

# Generative AI as a Pillar for Predicting 2D and 3D Wildfire Spread: Beyond Physics-Based Models and Traditional Deep Learning

@@ Abstract
Wildfires increasingly threaten human life, ecosystems, and infrastructure, with events like the 2025 Palisades and Eaton fires in Los Angeles County underscoring the urgent need for more advanced prediction frameworks. Existing physics-based and deep learning models struggle to capture dynamic wildfire spread across both 2D and 3D domains, especially when incorporating real-time, multimodal geospatial data. This paper explores how generative Artificial Intelligence (AI) models—such as GANs, VAEs, and Transformers—can serve as transformative tools for wildfire prediction and simulation. These models offer superior capabilities in managing uncertainty, integrating multimodal inputs, and generating realistic, scalable wildfire scenarios. We introduce a new paradigm that leverages large language models (LLMs) for literature synthesis, classification, and knowledge extraction, conducting a systematic review of recent studies applying generative AI to fire prediction and monitoring. We highlight how generative approaches uniquely address challenges faced by traditional simulation and deep learning methods. Finally, we outline five key future directions for generative AI in wildfire management, including unified multimodal modeling of 2D and 3D dynamics, agentic AI systems and chatbots for decision intelligence, and real-time scenario generation on mobile devices, along with a discussion of critical challenges. Our findings advocate for a paradigm shift toward multimodal generative frameworks to support proactive, data-informed wildfire response.
@@ 1 Introduction
Wildfires and bushfires have emerged as one of the most destructive natural disasters of the 21st century, leaving a devastating trail across natural ecosystems, agricultural lands, and densely populated urban regions Bowman et al. (2020); Doerr and Sant´ın (2016); Kotozaki (2025). These fast-moving fires not only cause immense structural damage and loss of human life but also result in widespread economic disruption and long-term environmental degradation Goralnick et al. (2025); Amiri et al. (2025); NASA Scientific Visualization Studio (2025). The release of smoke, fine particulate matter, and toxic gases during large-scale bushfires contributes significantly to air pollution, impacting public health and exacerbating climate change in cities across the globe Johnston et al. (2012); Urbanski (2014). Striking examples of this escalating crisis include the 2025 Los Angeles wildfire season—when the Palisades and Eaton fires swept through Southern California—and Australia’s 2019–2020 “Black Summer” fires Brown and Shelton (2025); Han et al. (2025); Woolcott (2025); Natural Hazards Research Australia (2023). Together, they caused an estimated \$250 billion in economic losses, destroyed thousands of homes, and displaced entire communities, inflicting deep physical and emotional suffering on affected populations (EPA, 2025; Gazette, 2025). These events underscore the urgent need for accurate and timely wildfire forecasting systems. In particular, real-time bushfire propagation simulation-capable of predicting fire spread pathways and identifying at-risk zones-plays a critical role in supporting firefighting operations, emergency evacuation planning, and long-term fire hazard mitigation strategies Baptiste Filippi et al. (2009); Barton et al. (2024); Wang and Zlatanova (2019); Wang et al. (2014).
@@
A variety of wildfire propagation models have been developed, each rooted in distinct computational approaches. Physics-based models (e.g., FARSITE, SPARK, Prometheus) simulate fire spread using physical laws of combustion, heat transfer, fuel conditions, and wind dynamics Finney (1998); Mell et al. (2007); Miller et al. (2015); Tymstra et al. (2010); Barton et al. (2020), offering detailed outputs but requiring extensive computation and environmental inputs—often limiting their use in real-time scenarios Dipierro et al. (2024). Empirical models, like McArthur’s Fire Danger Index, rely on historical fire behavior to produce rapid, region-specific predictions, but struggle with generalizability across diverse ecosystems Noble et al. (1980). Traditional machine learning models (e.g., decision trees, support vector machines) have been used for fire ignition and risk classification Jain et al. (2020); Bot and Borges (2022); Nur et al. (2023); Ismail et al. (2024), though their capacity to simulate dynamic fire progression is limited.
In contrast, deep learning models have recently gained momentum for 2D fire spread forecasting, due to their ability to capture complex spatiotemporal relationships in fire behavior Andrianarivony and Akhloufi (2024); Shadrin et al. (2024). Architectures such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), and U-Net variants have been applied to wildfire segmentation and short-term spread prediction, leveraging remote sensing, meteorological, and topographic data Ghali and Akhloufi (2023); Vargas (2025); Zakari et al. (2025); Alizadeh et al. (2024). These deep learning models (e.g., CNNs and RNNs) often outperform traditional physics-based simulators in runtime, scalability, and—in data-rich environments—predictive accuracy, as they learn complex patterns and variability directly from empirical observations rather than relying on explicit physical formulations Radke et al. (2019); Chatterjee et al. (2024). However, deep learning-based fire models still face key limitations that prevent them from enabling more advanced fire prediction capabilities with higher spatial and temporal resolution in real time. Most models remain constrained to 2D spatial domains Andrianarivony and Akhloufi (2024) and lack integration and real-time augmentation of multimodal inputs, such as meteorological, topographic, and vegetation (fuel) data Radke et al. (2019); Abdollahi and Yebra (2025), which often change dynamically as fire propagates across the landscape. In addition, many existing models cannot simulate vertical fire dynamics, limiting their effectiveness during fast-evolving wildfire events Xu et al. (2024d); He (2024).
Building on this trajectory, emerging generative AI models—such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and Transformers—offer substantial advantages over earlier deep learning architectures like CNNs and RNNs, particularly for complex environmental and urban modeling tasks Bhatia et al. ; Zhang and Wang (2024); Lifelo et al. (2024). These models already power a wide range of applications in everyday software for personal use, content creation, and entertainment, including image synthesis, text generation, music composition, and interactive conversational systems, where they deliver capabilities far beyond those of conventional deep learning approaches Li et al. (2025a). Unlike traditional models that primarily specialize in classification or regression, generative AI systems learn intricate data distributions to produce entirely new, coherent, and contextually appropriate outputs Sordo et al. (2025). This demonstrated ability to generate high-fidelity, multimodal content has sparked growing interest in adapting generative techniques to scientific and environmental domains, where similar benefits—such as robust uncertainty handling, scenario simulation, and the integration of heterogeneous data—hold significant transformative potential.
Motivated by this potential, this paper investigates how the convergence of generative AI techniques can drive next-generation bushfire prediction systems. We begin by conducting a systematic review of recent studies that apply generative AI to fire prediction and monitoring, utilizing a new paradigm that leverages large language models (LLMs) for literature synthesis, classification, and knowledge extraction. Through this review, we examine how these technologies support high-fidelity spatial modeling and improved scalability and accuracy, capabilities that are critical for enhancing situational awareness and decision-making during fire emergencies Dilo and Zlatanova (2011); Ma et al. (2024). We then outline five key future directions for advancing generative AI in wildfire management: developing unified multimodal frameworks that seamlessly integrate 2D and 3D data; designing conversational agentic AI systems for interactive, real-time wildfire intelligence; training interdisciplinary AI foundation models; enabling edge-based scenario generation on mobile and IoT devices; and advancing explainable AI interfaces to improve transparency and trust, accompanied by a discussion of critical challenges.
@@ 2 Previous Reviews in Fire Spread Management
Several previous studies have conducted comprehensive literature reviews on predicting bushfire spread using traditional methods, including simulation techniques, machine learning, and deep learning. While these approaches are not the primary focus of this paper—which centers on generative AI applications—they provide essential background on conventional fire spread prediction and its underlying scientific rationale. This section begins by summarizing these reviews, highlighting how both simulation models and deep learning approaches have been applied to fire propagation prediction. The overview of previous review articles aims to provide a comprehensive understanding of the existing AI applications in wildfire spread prediction and to identify gaps where generative AI models can be leveraged to address challenges faced by traditional simulation and deep learning approaches.
@@ 2.1 Wildfire Simulation Models and Traditional Machine Learning
Wildfire simulation models are traditional, well-established tools that simulate and predict bushfire spread by solving physical and empirical equations based on fuel, weather, and terrain conditions—originating from foundational models like Rothermel’s in the 1970s Rothermel (1972); Finney et al. (2012), and they have long served as the backbone of operational fire management and planning systems. Based on their underlying rationale, these models can be classified into several categories, as presented in Table 1, and are summarized through multiple previous studies Sullivan (2009); Finney et al. (2012).
@@
On the application level, a recent study conducts a comprehensive review of empirical and dynamic wildfire simulation models, focusing on their applications in predicting bushfire and wildfire spread across Australia Singh et al. (2025). It critically evaluates a suite of simulation systems including PHOENIX Rapidfire, SPARK, AUSTRALIS, REDEYE, IGNITE, and SiroFire, each leveraging distinct modeling techniques to simulate fire behavior, together with a variety of traditional machine learning and deep learning methods for wildfire prediction. PHOENIX Rapidfire, a deterministic simulator, uses a fire characterization model based on Huygen’s principle to predict spread, flame height, and intensity. While it excels in speed and ease of use, it is limited by reliance on predefined behavior models and underestimation of irregular fire shapes Pugnet et al. (2013); Richards (1995). SPARK provides flexible, modular simulation driven by user-defined spread models and integrates well with GIS platforms, though its high computational demands pose practical constraints. AUSTRALIS employs a discrete-event simulation based on empirically derived rate-of-spread models; despite its efficiency, it underperforms in severe fire conditions Miller et al. (2015). REDEYE and IGNITE integrate geospatial and real-time data for risk prediction and hotspot mapping, respectively, yet their effectiveness is limited by data compatibility and platform-specific requirements. SiroFire models dynamic weather components and supports strategic planning but lacks adaptability for diverse fuel types Singh et al. (2025). Collectively, these simulators provide valuable predictive capabilities; however, each exhibits distinct limitations, ranging from computational inefficiencies and limited model adaptability to reduced accuracy under extreme conditions. This review highlights the need for hybrid modeling approaches that integrate traditional simulation techniques with machine learning and real-time data to enhance predictive accuracy, reduce simulation runtime, and strengthen operational resilience.
@@ 2.2 Deep Learning in Wildfire Prediction
In the following subsection, we leverage a LLM to generate a literature summary and bibliographic visualizations that highlight the evolving trends in deep learning applications over time. Jain et al. (2020) presents a comprehensive scoping review of machine learning (ML) applications in wildfire science and management, analyzing 300 studies published up to the end of 2019. ML usage is categorized across six core domains: fuel characterization, fire detection, climate interactions, risk prediction, fire behavior, and post-fire effects. Among modern approaches, deep learning (DL) algorithms have shown strong potential for modeling wildfire spread, particularly due to their capacity to handle high-dimensional and multivariate data. CNNs are widely used for spatial fire detection and smoke recognition from imagery, while Long Short-Term Memory (LSTM) networks incorporate temporal dynamics to model fire growth. Deep Neural Networks (DNNs) have also been applied to map burned areas and forecast fire spread, often outperforming traditional models when sufficient annotated data are available. Despite their promise, the review emphasizes the importance of domain expertise and high-quality data for meaningful DL integration in operational forecasting.
A recent study systematically reviews deep learning (DL) applications in forest fire prediction, analyzing 55 key publications from 2017 to 2024 Mambile et al. (2024). The review emphasizes the transformative role of deep learning in capturing complex spatiotemporal patterns associated with fire ignition and spread, offering significant advantages over traditional machine learning approaches. CNNs are effectively used for analyzing satellite imagery to detect fire-prone areas and assess post-fire damage. LSTM and Gated Recurrent Unit (GRU) networks model time-series data to forecast fire progression based on historical environmental conditions. Emerging generative AI models such as GANs have been used for synthetic data generation, while Multilayer Perceptrons (MLPs) and U-Net architectures support fire risk estimation and boundary segmentation. Despite promising results, DL models face challenges such as data heterogeneity, limited inclusion of human activity factors, generalization across geographies, and the scarcity of high-quality labeled datasets. These limitations highlight the need for integrating diverse data sources and establishing standardized evaluation protocols to ensure model reliability in real-world wildfire prediction scenarios.
@@ 2.3 Limitations of the Existing Deep Learning Applications in Wildfire Prediction
Traditional models such as CNNs and RNNs typically produce deterministic outputs and struggle to quantify prediction uncertainty, an essential capability for wildfire management and emergency response applications involving stochastic environmental processes, such as abrupt changes in wind speed and directions Gal and Ghahramani (2016); Lakshminarayanan et al. (2017).
@@
Recurrent Neural Networks (RNNs) and Deep Neural Networks (DNNs) often struggle with vanishing gradients and limited temporal memory, which undermines their ability to capture long-range dependencies. This limitation makes them less effective in modeling the temporal progression of wildfires and in representing the long-term variability of underlying environmental processes such as fuel accumulation, climate patterns, and vegetation dynamics Hochreiter et al. (2001); Chen et al. (2024).
@@
Traditional deep learning architectures are not inherently designed to fuse multimodal data sources, such as 2D GIS data (e.g., satellite imagery, fuel load maps, and meteorological data) and 3D point clouds (e.g., digital terrain models) Li et al. (2022). As a result, generating multimodal fire spread prediction outputs across different dimensions—from 1D time series to 3D spatial representations—within a unified deep learning framework remains a significant challenge.
@@
Traditional deep learning models, such as CNNs and RNNs, typically achieve strong performance only when trained on abundant, high-quality labeled data. They depend heavily on large volumes of annotated samples, which are often scarce in wildfire prediction tasks due to the rare, unpredictable, and spatially heterogeneous nature of fire events Ghali and Akhloufi (2023); Xu et al. (2024d). However, many of these models lack the capability to perform data augmentation or generate synthetic training samples, limiting their predictive accuracy and generalizability in data-sparse or unseen regions. This limitation poses a significant challenge for developing robust wildfire.
@@
Environmental datasets frequently contain missing or incomplete information caused by sensor malfunctions, occlusion from cloud cover in satellite imagery, or data transmission failures Shen et al. (2015). Such data gaps hinder accurate modeling and prediction. Traditional deep learning models often assume complete input data or rely on simplistic imputation methods that fail to capture the underlying spatiotemporal dependencies critical for wildfire dynamics. These limitations reduce model robustness and prediction accuracy in real-world scenarios where data is inherently noisy or sparse Li et al. (2020).
@@
Deep learning models such as CNNs and RNNs often operate as ”black boxes,” providing limited insight into how predictions are made and offering little transparency or trustworthiness Rudin (2019); Doshi-Velez and Kim (2017).